satellite73 (satellite73):

what math teacher made up this question? Glenn needs 20 minutes to clean his room. If his little sister Veronika pulls things out while he is cleaning, it takes 60 minutes for him to clean the room. How many minutes does it take for Veronika to pull everything out?

OpenStudy (empty):

I feel like the answer I want to give is not what they're thinking the answer _should_ be.

OpenStudy (anonymous):

my answer is five seconds , but the answer they want is 30 minutes has to be one of the dumbest questions i have seen lately !

OpenStudy (empty):

I think the solution is simple: waterboarding.

OpenStudy (anonymous):

lol

OpenStudy (usukidoll):

why do they want 30 minutes though? just get a vacuum cleaner and get rid of all the junk in the room. Takes 5 seconds XD

OpenStudy (anonymous):

i know it is 30 minutes because of the answer choices given they want the solution to \[\frac{1}{20}-\frac{1}{x}=\frac{1}{60}\] as if that has anything to do with it right to the "bad math" pile

OpenStudy (astrophysics):

I remember you asking another question similar to this, it was the exact same solution

OpenStudy (anonymous):

must be a franchise franchise of math education majors

OpenStudy (empty):

A franchise destined for Guantanamo Bay?

OpenStudy (astrophysics):

lol

OpenStudy (usukidoll):

I didn't write those problems though lol. it's probably some mad person who wrote it and want to torture people x.x !

OpenStudy (astrophysics):

I don't really understand this solution

OpenStudy (astrophysics):

why 1/20-1/x=1/60

OpenStudy (usukidoll):

 gonna try to attempt this wazoo.